OpenStudy (anonymous):

please help!!! solve the system by elimination y=x^2+2 -4x-y=10 Part a: solve the following sytem of equations. round to the nearest hundredth if necessary. y=-2x+7 y= -x^2-4x+7 Part B: graph the system

OpenStudy (dumbcow):

are these 2 different systems ?

OpenStudy (dumbcow):

with a quadratic equation in the system... i would recommend using substitution, not elimination

OpenStudy (radar):

Write each equation in terms of y y=x^2+2 y=-4x-10 y=y x^2+2=-4x-10 then we would have x^2+4x+12=0 A quadratic equation which can be solved using the quadratic formula.

OpenStudy (anonymous):

so how would i use substitution?

OpenStudy (radar):

We have already eliminated y, solve for x.

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

how do i solve for x? sorry

OpenStudy (radar):

\[-b \pm \sqrt{b ^{2}-4ac}\over 2a\]a=1 b=4 c=12

OpenStudy (radar):

plug you values in the equation and solve for x, after obtaining the x values then solve for y using the first equation.

OpenStudy (radar):